# Tracing attosecond currents and controlling their direction in a scanning tunneling microscope

### v1

@@
Abstract
The tunneling effect driven by waveform-controlled laser pulses is a cornerstone of attosecond science and contributes decisively to its extreme time resolution. In the spatial domain, electron tunneling through a bias-driven junction between a nanotip and a surface enables atomic-scale spatial resolution in scanning tunneling microscopy (STM). Recently, first signatures of attosecond modulation of STM currents have shown that simultaneous attosecond-nanometer resolution is feasible. However, while sub-cycle attosecond dynamics are routinely controlled and determined with high precision, controlling the direction of the currents and determining their duration have remained elusive in STM. Here, we induce STM tunneling currents using two-color laser pulses and dynamically control their direction, relying solely on the sub-cycle waveform of the pulses. Projecting our measurement data onto numerical and analytical solutions of the time-dependent Schrödinger equation reveals non-adiabatic tunneling as the underlying physical mechanism, yielding current burst durations on the order of 900 as. Despite working under ambient conditions but free from thermal artifacts, we achieve a lateral spatial resolution of 2 nm and sub-angström topographic sensitivity. Directional control of attosecond bursts in STM will enable triggering and detecting ultrafast charge dynamics in molecular systems and defect states at the spatio-temporal frontier of microscopy.
@@ Introduction
Attosecond science is based on the control of electron motion on the atomic time scale by strong electric fields 1. While it has enabled the observation of a wide range of attosecond phenomena, achieving a spatial resolution on the order of the atomic length scale in the angström-nanometer regime is challenging. The main pathway to attosecond microscopy is the use of attosecond-scale electron pulses instead of light due to their small de Broglie wavelength, typically much less than 1 nm. In addition, electrons must be spatially confined, either as a recolliding electron wavepacket originating from an atom, molecule, or nanostructure 2, 3, 4, 5, as a narrow beam in an electron microscope 6, 7, or inside an atomic-scale tunneling microscopy junction 8, 9, 10. The strength of the latter approach is that scanning tunneling microscopy (STM) in its static implementation using a bias field provides atomic resolution and precise energy resolution out of the box for a wide range of systems, such as molecules 11. STM tunneling currents driven by single-cycle THz transients 12, which can be regarded as a slowly varying static bias field, have enabled real-time observations of atomic-scale molecular vibrations on the sub-picosecond time scale 13 and exciton formation in molecules 14, for instance.
@@
Tuning the carrier-envelope phase (CEP) can shape an ultrashort THz field waveform to break its symmetry and obtain a field that is much stronger in one direction compared to the other. Maximum tunneling is then achieved by the field in that direction, enabling control of the dominant direction of the current. This allows, for example, the controlled extraction of an electron from the sample 15, an important prerequisite for pump-probe measurements 13. Moving to infrared wavelengths and the femtosecond and attosecond time scales, CEP modulation is a well-known approach to control the sub-cycle electron motion in atoms 16, 17, 18, metallic nanotips 19, 20, 5, nanodevices 21 and vacuum nanogaps 22, 23, 24 and enables attosecond time resolution in these systems. It may seem straightforward to induce attosecond tunneling currents and control their direction in STM using CEP-stable infrared laser pulses, building on many experiments exploring ultrafast STM in the femtosecond regime (see, e.g., 25, 26, 27, 28, 29, 30, 10). However, pioneering ultrafast STM experiments with infrared lasers have shown that thermal artifacts are often present when conventional lock-in approaches are used to extract low-level laser-induced currents 25, 31. An encouraging experiment has reported signatures of CEP-controlled tunneling current modulation in STM using 6-fs near-infrared laser pulses 8. Still, robust sub-cycle control of the direction of STM currents and timing of their dynamics have remained elusive.
@@
In this work, we demonstrate robust attosecond directional control of ultrafast tunneling currents in an STM junction. We employ two-color pulses, a powerful alternative to CEP-stable pulses, and achieve ultrafast measurements without thermal artifacts. To this end, we superimpose an infrared laser pulse with its second harmonic in the same polarization plane and change their relative time delay. Two-color pulses lead to field waveform with controlled symmetry breaking and enables sub-cycle control, as demonstrated in a wide range of attosecond phenomena 32, 33, 34, 35, 4. Using the synthesized field waveform of the two-color pulse, we are able to seamlessly switching between tip-to-sample tunneling and vice versa. From an excellent agreement of experiment and theory we are able to identify the underlying physical mechanism as non-adiabatic tunneling, with electron burst durations of as. Free of thermal artifacts, our microscope enables a lateral spatial resolution of 2 nm despite operation under ambient conditions.
@@ Results
Ultrafast two-color scanning tunneling microscopy
@@
In our experiment, we perform STM at ambient conditions at room temperature using two-color laser pulses (see Fig. 1a). An electropolished Pt:Ir nanotip with an apex radius of 5…20 nm and a gold substrate prepared by a self-formation approach 36 form a tunneling junction with variable width . We irradiate the junction with 35 fs, 1850 nm laser pulses of up to 75 pJ pulse energy at a repetition rate of 80 MHz together with its second harmonic (SH). The intensity ratio of the SH and the fundamental is 10% in order to induce a pronounced symmetry breaking of the resulting waveform (see the inset of Fig. 1a for an illustration). Figure 1b shows the optical setup of our experiment (see Methods for a detailed description). We control the relative time delay between the two colors using an interferometric setup with the help of a linear delay stage. The two-color pulses induce an optical near-field in the nano-scale tunneling junction, leading to a strong field enhancement and enabling us to enter the strong-field sub-cycle regime of light-matter interaction.
@@
The microscope is operated in constant-current measurement mode, where a static bias field drives a tunneling current and a feedback loop controls the relative distance between tip and sample to achieve a constant current 37. In addition, ultrafast laser-induced currents are generated by the optical near-field formed inside the tunneling junction through irradiation with the two-color laser pulses. Isolating the contribution of the laser-induced currents from the static tunneling current is a challenging task. A chopper-based lock-in approach (cf. 8) has not been applicable because it produces artifact signals. Periodic on-off switching of the intensity in the kHz domain leads to expansion and contraction of the nanotip and a corresponding increase and decrease of the static tunneling current, masking the actual laser-induced current signal (see Supplementary Information). In order to avoid any intensity modulation and the artifacts accompanying it, we modulate the two-color delay with a sinusoidal modulation 28 at an angular frequency of kHz using a mirror mounted on a vibrating piezoelectric chip. The total two-color delay as a function of time is then given by , where is the base delay set by the linear delay stage and is the amplitude of the sinusoidal delay modulation. This leads to a clear and consistent lock-in signal and allows us to measure the laser-induced current independently of the microscope feedback loop operating at a locking bandwidth below 500 Hz.
@@
Figure 2 shows a scan of the gold sample under laser irradiation for a fixed value of . The sample bias voltage is 200 mV and the set-point current for the conventional STM operation is 100 pA. The topography map (Fig. 2a) obtained from the static tunneling current reveals fine atomic steps of the gold sample on the angström scale (see also the line cross section in Fig. 2b). Simultaneously with the topography, the lock-in signal at 3.7 kHz allows us to record a map of the laser-induced current (Fig. 2c), which largely follows the topographic features of the sample. The phase offset of the laser-induced current signal with respect to the delay modulation (Fig. 2d) is found to be approximately uniform, demonstrating that we are indeed detecting a laser-induced signal that depends on the two-color delay with sub-cycle resolution. A closer look at the line cross-section (see Fig. 2b) shows that the laser-induced current signal is sensitive to the tiny variations of the topographic height of the sample on the sub-angström scale. We find that peaks in the signal are correlated with small topographic protrusions, hinting at a remarkable sensitivity of the optical near-field to the atomic-scale sample geometry in the STM junction 38, 39. We estimate the lateral spatial resolution obtained from the laser-induced signal to be around 2 nm, roughly the same as the spatial resolution of the conventional topographic scan. Here we are likely limited by the fact that the microscope is operated in ambient conditions.
Sub-cycle waveform control of tunneling currents
@@
Scanning the base two-color delay with a step size of 130 as and placing the nanotip at a fixed spot on the sample allows us to study the sub-cycle dependence of the laser-induced current. Figure 3a shows the lock-in current as a function of (see Supplementary Information for a plot of a wide-range delay scan). We find oscillations with a period of 1.5 fs, which corresponds to half of the period of the SH field, a clear sub-cycle feature. The lock-in measurement also yields the phase of the lock-in current signal (Fig. 3b). We find alternating phase jumps of 180 degrees on top of an arbitrary but constant phase offset. A procedure to reconstruct the actual laser-induced current from the lock-in signal (see Supplementary Information for details) reveals that controls the magnitude of the current, as shown by the dashed orange curve in Fig. 3c; it may also indicate control of the direction of the current. However, the reconstruction from the lock-in measurements alone cannot provide any information about the offset of the current which remains undefined. A different, more direct measurement of the laser-induced current is required.
@@
In order to show that indeed controls the direction of the laser-induced current, we temporarily switch off the microscope feedback loop as well as the lock-in modulation of the two-color delay (). We freeze the position of the nanotip in space and set the sample bias to zero, thus suppressing the static tunneling current and preempting any symmetry breaking by the static field. The only measurement quantity that remains is the net laser-induced current, now measured directly. The red curve in Fig. 3c shows the resulting signal, which we find to be oscillatory with the period of the SH field. The signal is in good agreement with the laser-induced current reconstructed from the lock-in approach. Our measurements unambiguously demonstrate that the direction of the current is governed by . Positive currents indicate that the asymmetric waveform of the two-color field drives electrons in sub-cycle bursts from the nanotip to the sample. For negative currents, the opposite is true – we obtain a current from the sample to the nanotip. Here, it is solely the laser field that breaks the symmetry of the tunneling junction and produces a net current, a hallmark of lightwave electronics, where the sub-cycle waveform of a laser pulse governs ultrafast electric current dynamics 40. Since our laser pulse comprises many cycles, we obtain a train of attosecond current bursts with a periodicity of the fundamental field.
@@
Figure 3d shows the polarization dependence of the laser-induced current measured at zero bias while the nanotip is frozen. The polarization angle is defined with respect to the sample plane. The laser-induced current peaks sharply at a polarization angle of 100 degrees, which coincides approximately with the scenario in which the two-color laser field is aligned with the perpendicularly oriented nanotip. This ensures maximum near-field enhancement in the gap and therefore maximum laser-induced current. At 280 degrees, the same scenario is reached, but the flip of the polarization causes a reversal of the sign of the two-color waveform. For the two angles, we perform a two-color delay scan and indeed observe the resulting reversal of the direction of the current (Fig. 3e). This further corroborates the notion that the waveform controls the transport of electrons through the junction.
Electron transport in the non-adiabatic tunneling regime
@@
A crucial question which needs to be answered is the underlying physics of the electron transport. In agreement with photoemission physics from freestanding nanotips 41, experiment 8 and theory 42, 43, 44 show that the dichotomy of a multiphoton regime and a laser-induced tunneling regime, parametrized by the Keldysh parameter 45 , also governs ultrafast STM. In STM, the only measurement observable is the average current, unlike in nanotip photoemission experiments, for instance, where the photoelectron spectrum can reveal crucial information about the underlying physics (see, e.g. 19, 20, 4). Here we measure the scaling of the laser-induced current with total laser power (SH and fundamental), shown in Fig. 4a in a double-logarithmic scale. Freezing the tip at zero bias, we determine the power scaling of the amplitude of the two-color current modulation for three different junction widths, free from any other signals. We observe that the current increases approximately like a power law with a slope of , which may indicate a multiphoton regime following the absorption of 3 or 4 photons from the two-color pulse. However, beyond a power of 5 mW, the nonlinearity begins to decrease. A soft kink appears in the scaling, a signature of the transition from the multiphoton to the tunneling regime of ultrafast STM 8, 42, in analogy to photoemission from nanostructures 41, 46. Another effect that may contribute to the kink is the fact that we begin to transfer more than one electron per laser pulse starting at an average current of 13 pA.
@@
In order to gain a deeper insight into the physical mechanism, we employ a simple theory model, a numerical integration of the time-dependent Schrödinger equation (nTDSE, see Methods and Supplementary Information for details) and directly compare the experiment with our calculations, taking into account all known experimental parameters (see Fig. 4a). Here, we account for the fact that the optical near-field enhancement, which translates the total laser power into the actual field driving electrons across the tunneling junction, decreases with increasing junction width. The calculated curves allow us to extract field enhancement factors for the fundamental field of , and for 1 nm, 1.5 nm and 2 nm, respectively. These values are qualitatively in agreement with a numerical solution of the Maxwell equations for the tip-sample geometry (see Supplementary Information). The excellent agreement between experiment and simple single-electron theory allows us to determine the effective Keldysh parameter (see the Supplementary Information for the derivation of for a two-color field). The upper horizontal axis in Fig. 4a shows for a junction width of 1 nm. We find that the dominating mechanism is non-adiabatic tunneling in the vicinity of . Here, an electron tunnels through the potential barrier strongly suppressed by the peak of the near-field (see Fig. 4b for an illustration for electron transport from the tip to the sample). However, due to the fact that the barrier cannot be considered as quasi-static in this regime and varies strongly during the evolution of the tunneling process, the electron is gaining energy “under the tunneling barrier” (cf. 47, 48). In our case, non-adiabatic tunneling is the first step of the three-step mechanism of ultrafast STM 42, 43, 44. In the second step, the electron exits the tunneling barrier, is further accelerated, and is transmitted into the adjacent metal in the third and final step.
@@
The sub-cycle waveform control of the direction of the attosecond current bursts can also be explained by the nTDSE model and by an analytical strong-field (SF) model based on the van Vleck propagator 44 (see Methods and Supplementary Information for details). Figure 5a shows the laser-induced current measured from Fig. 3c. In addition, we display the theory curves obtained from the nTDSE and SF models for a peak field strength of of the fundamental () after scaling their amplitude to match the experimental data. We find excellent agreement. The same type of sinusoidal oscillation is found in both theory models within the range of , corresponding to (see Supplementary Information). The experimentally observed current modulation is nearly perfectly symmetric with respect to zero current, which attests to the fact that the symmetry breaking induced by the two-color waveform dominates over all other symmetry-breaking factors, for example the stark difference in geometry between nanotip and sample.
@@
The excellent overall agreement between experiment and theory allows us to trace the electron dynamics and estimate the duration of the attosecond electron bursts. Figure 5b shows the current density map calculated with the nTDSE as a function of space and time across a 1 nm junction for a peak field strength of () and for maximum symmetry breaking. The electrons are created at the peak of the field (time zero) and cross the barrier within 1 fs without quiver motion 23, 42, 43. The full-width at half maximum (FWHM) duration of the electron burst at the center is shown in Fig. 5c and amounts to 840 as. This is shorter than 1.2 fs obtained in the single-color case (1850 nm field only), showing that the two-color approach can not only induce symmetry breaking, but also help reach the sub-femtosecond time domain. The flow of the current density matches well with semiclassical trajectories obtained from the SF model that appear after the tunneling exit located at around 0.48 nm (see dotted curves in Fig. 5b). Figure 5d shows the dependence of the induced symmetry breaking in the current flow on the intensity ratio of the SH field to the fundamental field, quantified by the directionality . Here, and are the integrated currents in positive and in negative direction, respectively. Starting at low intensity ratios, increases strongly and reaches a saturation regime around an intensity ratio of 0.15. At 8 V nm-1 and a ratio of 0.2, we reach , indicating that most electrons follow the preferred direction set by the field waveform.
Discussion

### v2

# Clocking and controlling attosecond currents in a scanning tunnelling microscope

@@
Clocking and controlling attosecond currents in a scanning tunnelling microscope
Quantum tunnelling of electrons can be confined to the sub-cycle time scale of strong light fields 1, 2, contributing decisively to the extreme time resolution of attosecond science 3. Because tunnelling also enables atomic-scale spatial resolution in scanning tunnelling microscopy (STM 4), integrating STM with light pulses has long been a key objective in ultrafast microscopy, spanning the picosecond 5, 6, 7, 8 and femtosecond domains 9, 10, 11, with first signatures of attosecond dynamics 12. However, while sub-cycle dynamics on the attosecond time scale are routinely controlled and determined with high precision 1, 13, 2, 14, 15, 16, 17, 18, 19, controlling the direction of attosecond currents and determining their duration have remained elusive in STM. Here, we induce STM tunnelling currents using two-colour laser pulses and dynamically control their direction, relying solely on the sub-cycle waveform of the pulses. Projecting our measurement data onto numerical and analytical solutions of the time-dependent Schrödinger equation reveals non-adiabatic tunnelling as the underlying physical mechanism, yielding a current burst duration of 860 as. Despite working under ambient conditions but free of thermal artifacts, we achieve sub-angström topographic sensitivity and a lateral spatial resolution of 2 nm. This unprecedented capability to directionally control attosecond bursts will enable triggering and imaging ultrafast charge dynamics in atomic, molecular and condensed systems at the spatio-temporal microscopy frontier of lightwave electronics 20, 21, 22.
@@ Introduction
Attosecond science is based on the control of electron motion on the atomic time scale by strong electric fields 3. While it has enabled the observation of a wide range of attosecond phenomena, achieving a spatial resolution on the order of the atomic length scale in the angström-nanometer regime is challenging. The main pathway to attosecond microscopy is the use of attosecond-scale electron pulses instead of light due to their small de Broglie wavelength, typically much less than 1 nm. In addition, electrons must be spatially confined, either as a recolliding electron wavepacket originating from an atom, molecule, or nanostructure 23, 24, 19, 18, as a narrow beam in an electron microscope 25, 26, or inside an atomic-scale tunnelling microscopy (STM) junction 12, 27, 11. The strength of the latter approach is that STM in its static implementation using a bias field provides atomic resolution and precise energy resolution out of the box for a wide range of systems, such as molecules 28. STM tunnelling currents driven by single-cycle THz transients 5, which can be regarded as a slowly varying static bias field, have enabled real-time observations of atomic-scale molecular vibrations on the sub-picosecond time scale 6 and exciton formation in molecules 8, for instance.
@@
Tuning the carrier-envelope phase (CEP) can shape an ultrashort THz field waveform to break its symmetry and obtain a field that is much stronger in one direction compared to the other. Maximum tunnelling is then achieved by the field in that direction, enabling control of the dominant direction of the current. This allows, for example, the controlled extraction of an electron from the sample 7, an important prerequisite for pump-probe measurements 6. Moving to infrared wavelengths and the femtosecond and attosecond time scales, CEP modulation is a well-known approach to control the sub-cycle electron motion in atoms 29, 30, 1, metallic nanotips 13, 31, 18, nanodevices 17 and vacuum nanogaps 15, 16, 32 and enables attosecond time resolution in these systems. It may seem straightforward to induce attosecond tunnelling currents and control their direction in STM using CEP-stable infrared laser pulses, building on many experiments exploring ultrafast STM in the femtosecond regime (see, e.g., 9, 33, 10, 34, 35, 36, 11). However, pioneering ultrafast STM experiments with infrared lasers have shown that thermal artifacts are often present when conventional lock-in approaches are used to extract low-level laser-induced currents 9, 37. First signatures of CEP-controlled tunnelling current modulation in STM using 6-fs near-infrared laser pulses were reported recently 12. Still, sub-cycle control of the direction of STM currents and timing of their dynamics have remained elusive. This advance is necessary for attosecond imaging of ultrafast electron phenomena.
@@
In this work, we demonstrate robust attosecond directional control of ultrafast tunnelling currents in an STM junction. We employ two-colour pulses, a powerful alternative to CEP-stable pulses, and achieve ultrafast measurements without thermal artifacts. To this end, we superimpose an infrared laser pulse with its second harmonic in the same polarization plane and change their relative time delay. Two-colour pulses lead to field waveform with controlled symmetry breaking and enables sub-cycle control, as demonstrated in a wide range of attosecond phenomena 38, 39, 40, 41, 19. Using the synthesized field waveform of the two-colour pulse, we are able to seamlessly switch between tip-to-sample tunnelling and vice versa. From an excellent agreement of experiment and theory we are able to identify the underlying physical mechanism as non-adiabatic tunnelling, with electron burst durations of as. Free of thermal artifacts, our microscope enables a lateral spatial resolution of 2 nm despite operation under ambient conditions.
@@ Results
Ultrafast two-colour scanning tunnelling microscopy
@@
In our experiment, we perform STM at ambient conditions at room temperature using two-colour laser pulses (see Fig. 1a). An electropolished Pt:Ir nanotip with an apex radius of 5…20 nm and a gold substrate form a tunnelling junction with variable width . We irradiate the junction with 35 fs, 1850 nm laser pulses of up to 75 pJ pulse energy at a repetition rate of 80 MHz together with its second harmonic (SH). The intensity ratio of the SH and the fundamental is 10% in order to induce a pronounced symmetry breaking of the resulting waveform (see the inset of Fig. 1a for an illustration). Figure 1b shows the optical setup of our experiment (see Methods for a detailed description). We control the relative time delay between the two colours using an interferometric setup with the help of a linear delay stage. The two-colour pulses induce an optical near-field in the nano-scale tunnelling junction, leading to a strong field enhancement and enabling us to enter the strong-field sub-cycle regime of light-matter interaction.
@@
The microscope is operated in constant-current measurement mode, where a static bias field drives a tunnelling current and a feedback loop controls the relative distance between tip and sample to achieve a constant current. In addition, ultrafast laser-induced currents are generated by the optical near-field formed inside the tunnelling junction through irradiation with the two-colour laser pulses. Isolating the contribution of the laser-induced currents from the static tunnelling current is a challenging task. A chopper-based lock-in approach (cf. 12) has not been applicable because it produces artifact signals. Periodic on-off switching of the intensity in the kHz domain leads to expansion and contraction of the nanotip and a corresponding increase and decrease of the static tunnelling current, masking the actual laser-induced current signal (see Supplementary Information). In order to avoid any intensity modulation and the artifacts accompanying it, we modulate the two-colour delay with a sinusoidal modulation 34 at an angular frequency of kHz using a mirror mounted on a vibrating piezoelectric chip. The total two-colour delay as a function of time is then given by , where is the base delay set by the linear delay stage and is the amplitude of the sinusoidal delay modulation. This leads to a clear and consistent lock-in signal and allows us to measure the laser-induced current independently of the microscope feedback loop operating at a locking bandwidth below 500 Hz.
@@
Figure 1c-e shows a scan of the gold sample under laser irradiation for a fixed value of . The sample bias voltage is 200 mV and the set-point current for the conventional STM operation is 100 pA. The topography map (Fig. 1a) obtained from the static tunnelling current reveals fine atomic steps of the gold sample on the angström scale. Simultaneously with the topography, the lock-in signal at 3.7 kHz allows us to record a map of the laser-induced current (Fig. 1d), which largely follows the topographic features of the sample. The phase offset of the laser-induced current signal with respect to the delay modulation is found to be approximately uniform (see Supplementary Information). A look at a line cross-section (see Fig. 1d) shows that the laser-induced current signal is sensitive to the tiny variations of the topographic height of the sample on the sub-angström scale. We find that peaks in the signal are correlated with small topographic protrusions, hinting at a remarkable sensitivity of the optical near-field to the atomic-scale sample geometry at the STM junction 42, 43. We estimate the lateral spatial resolution obtained from the laser-induced signal to be around 2 nm, roughly the same as the spatial resolution of the conventional topographic scan. Here we are likely limited by the fact that the microscope is operated in ambient conditions.
Sub-cycle waveform control of tunnelling currents
@@
Scanning the base two-colour delay with a step size of 130 as and placing the nanotip at a fixed spot on the sample allows us to study the sub-cycle dependence of the laser-induced current. Figure 2a shows the lock-in current as a function of (see Supplementary Information for a plot of a wide-range delay scan). We find oscillations with a period of 1.5 fs, which corresponds to half of the period of the SH field, a clear sub-cycle feature. The lock-in measurement also yields the phase of the lock-in current signal (Fig. 2b). We find alternating phase jumps of 180 degrees on top of an arbitrary but constant phase offset. A procedure to reconstruct the actual laser-induced current from the lock-in signal (see Supplementary Information for details) reveals that controls the magnitude of the current, as shown by the dashed orange curve in Fig. 2c; it may also indicate control of the direction of the current. However, the reconstruction from the lock-in measurements alone cannot provide any information about the offset of the current which remains undefined. A different, more direct measurement of the laser-induced current is required.
@@
In order to show that indeed controls the direction of the laser-induced current, we temporarily switch off the microscope feedback loop as well as the lock-in modulation of the two-colour delay (). We freeze the position of the nanotip in space and set the sample bias to zero, thus suppressing the static tunnelling current and preempting any symmetry breaking by the static field. The only measurement quantity that remains is the net laser-induced current, now measured directly. The red curve in Fig. 2c shows the resulting signal, which we find to be oscillatory with the period of the SH field. The signal is in good agreement with the laser-induced current reconstructed from the lock-in approach. Our measurements unambiguously demonstrate that the direction of the current is governed by . Positive currents indicate that the asymmetric waveform of the two-colour field drives electrons in sub-cycle bursts from the nanotip to the sample. For negative currents, the opposite is true – we obtain a current from the sample to the nanotip. Here, it is solely the laser field that breaks the symmetry of the tunnelling junction and produces a net current, a hallmark of lightwave electronics, where the sub-cycle waveform of a laser pulse governs ultrafast electric current dynamics 21. Since our laser pulse comprises many cycles, we obtain a train of attosecond current bursts with a periodicity of the fundamental field.
@@
Figure 2d shows the polarization dependence of the laser-induced current measured at zero bias while the nanotip is frozen. The polarization angle is defined with respect to the sample plane. The laser-induced current peaks sharply at a polarization angle of 100 degrees, which coincides approximately with the scenario in which the two-colour laser field is aligned with the perpendicularly oriented nanotip. This ensures maximum near-field enhancement in the gap and therefore maximum laser-induced current. At 280 degrees, the same scenario is reached, but the flip of the polarization causes a reversal of the sign of the two-colour waveform. For the two angles, we perform a two-colour delay scan and indeed observe the resulting reversal of the direction of the current (Fig. 2e). This further corroborates the notion that the waveform controls the transport of electrons through the junction.
Electron transport in the non-adiabatic tunnelling regime
@@
A crucial question which needs to be answered is the underlying physics of the electron transport. An ultrafast STM experiment 12 and theory studies 44, 45, 46 show that the dichotomy of a multiphoton regime and a laser-induced tunnelling regime, parametrized by the Keldysh parameter 47 , also governs ultrafast STM. In STM, the only measurement observable is the average current, unlike in nanotip photoemission experiments, for instance, where the photoelectron spectrum can reveal crucial information about the underlying physics (see, e.g. 13, 31, 19). Here we measure the scaling of the laser-induced current with total laser power (SH and fundamental), shown in Fig. 3a in a double-logarithmic scale. Freezing the tip at zero bias, we determine the power scaling of the amplitude of the two-colour current modulation for three different junction widths, free from any other signals. We observe that the current increases approximately like a power law with a slope of , which may indicate the absorption of 3 or 4 photons from the two-colour pulse. However, beyond a power of 4 mW, the nonlinearity begins to decrease. A soft kink appears in the scaling, a signature of the transition to the tunnelling regime of ultrafast STM 12, 44, analogous to photoemission from nanostructures 48, 49. Another effect that may contribute to the kink is the fact that we begin to transfer more than one electron per laser pulse starting at an average current of 13 pA.
@@
In order to gain a deeper insight into the physical mechanism, we employ a simple theory model, a numerical integration of the time-dependent Schrödinger equation (nTDSE, see Supplementary Information for details) and directly compare the experiment with our calculations, taking into account all known experimental parameters (see Fig. 4a). Here, we account for the fact that the optical near-field enhancement, which translates the total laser power into the actual field driving electrons across the tunnelling junction, decreases with increasing junction width. The calculated curves allow us to extract field enhancement factors for the fundamental field of , and for 1 nm, 1.5 nm and 2 nm, respectively. These values are qualitatively in agreement with numerical solutions of the Maxwell equations for the tip-sample geometry (see Supplementary Information). The excellent agreement between experiment and simple single-electron theory allows us to determine the effective Keldysh parameter (see the Supplementary Information for the derivation of for a two-colour field). The upper horizontal axis in Fig. 3a shows for a junction width of 1 nm. We find that the dominating mechanism is non-adiabatic tunnelling in the vicinity of . Here, an electron tunnels through the potential barrier strongly suppressed by the peak of the near-field (see Fig. 3b for an illustration for electron transport from the tip to the sample). However, due to the fact that the barrier cannot be considered as quasi-static in this regime and varies strongly during the evolution of the tunnelling process, the electron is gaining energy “under the tunnelling barrier” by photon absorption (cf. 50, 51, 46). In our case, non-adiabatic tunnelling is the first step of the three-step mechanism of ultrafast STM 44, 45, 46. In the second step, the electron exits the tunnelling barrier, is further accelerated, and is transmitted into the adjacent metal in the third and final step.
@@
The sub-cycle waveform control of the direction of the attosecond current bursts can also be explained by the nTDSE model and by an analytical strong-field (SF) model based on the van Vleck propagator 46 (see Methods and Supplementary Information for details). Figure 4a shows the laser-induced current measured from Fig. 2c. In addition, we display the theory curves obtained from the nTDSE and SF models for a peak field strength of of the fundamental () after scaling their amplitude to match the experimental data. We find excellent agreement. The experimentally observed current modulation is nearly perfectly symmetric with respect to zero current, which attests to the fact that the symmetry breaking induced by the two-colour waveform dominates over all other symmetry-breaking factors, for example the stark difference in geometry between nanotip and sample.
@@
The excellent overall agreement between experiment and theory allows us to trace the electron dynamics and estimate the duration of the attosecond electron bursts. Figure 4b shows the current density map calculated with the nTDSE as a function of space and time across a 1 nm junction for a peak field strength of () and for maximum symmetry breaking. The electrons are created at the peak of the field (time zero) and cross the barrier on a scale much shorter than the optical cycle duration and hence without any quiver motion 16, 44, 45. The flow of the current density matches well with semiclassical trajectories obtained from the SF model that appear after the tunnelling exit located at 0.35 nm (see dotted curves in Fig. 4b). The full-width at half maximum (FWHM) duration of the electron burst at the end of the junction at nm is depicted in Fig. 4c and amounts to 860 as. This is shorter than 970 as obtained in the single-colour case (1850 nm field only), showing that the two-colour approach not only induces full symmetry breaking, but also leads to a shorter burst duration within the sub-femtosecond time domain. In Fig. 4c, we can also observe that the arrival time is about 700 as after the peak of the field, which accounts for the travel time through the junction. The burst duration is strongly confined to the sub-cycle regime for a wider range of peak field strengths (see Fig. 4d), showing that attosecond current bursts can be generated routinely. The same is true for variations of more experimental parameters, such as the intensity ratio of the SH and the fundamental, junction width and tip workfunction; the burst duration stays well below the 1 fs mark (see Supplementary Information). While our experiment is performed at an intensity ratio of the SH and the fundamental of 10%, ratios as small as 2.5% can induce strong symmetry breaking (see Supplementary Information).
Conclusion